On a spelling test, Mariah had 1/6 of her words wrong. If 15 words were wrong, how many were right?
Answer by stanbon(75887)

On a spelling test, Mariah had 1/6 of her words wrong. If 15 words were wrong, how many were right?
---------------------
If 1/6 were wrong, 5/6 were correct.
---
Equation:
Let "x" be be the total number in the test.
(1/6)x = 15
x = 90
---
# correct = (5/6)90 = 75
================================
Cheers,
Stan H.
================================